### v2

# Leveraging Soft Prompts for Privacy Attacks in Federated Prompt Tuning

Abstract
Membership inference attacks (MIAs) pose a serious privacy threat in federated learning (FL). While MIAs have been extensively studied in standard FL, the recent shift toward federated fine-tuning introduces new and largely unexplored attack surfaces. In this work, we show that federated prompt-tuning, which adapts pre-trained models using lightweight input prefixes, exposes a novel and effective vector for membership inference. We propose PromptMIA, a membership inference attack tailored to federated prompt-tuning, in which a malicious server introduces adversarially crafted prompts and exploits their updates during collaborative training to determine whether a target data point belongs to a client’s private dataset. We formalize this threat via a security game and demonstrate that PromptMIA achieves consistently high attack advantage across diverse benchmark datasets. We also provide a theoretical lower bound on the attack advantage that explains the observed empirical behavior. Finally, we show that existing MIA defenses are often ineffective against PromptMIA, highlighting the need for defense mechanisms specifically tailored to prompt-tuning in federated settings.
Machine Learning, ICML
1 Introduction
Federated Learning (FL) (McMahan et al., 2017) is a prominent model training paradigm that enables data holders to collaboratively train a shared model without exposing their private data. However, while FL methods are privacy-compliant, they remain vulnerable to adversarial threats (Muñoz-González et al., 2017; Zhu et al., 2019). One prominent threat is membership inference attack (MIA)  (Shokri et al., 2017), where the adversarial server attempts to determine whether a specific data record was included in a client’s training dataset. Existing research has predominantly studied such privacy threats in settings where the attack surface is tied to clients’ gradients or full model parameters exchanged during training (Melis et al., 2019; Li et al., 2023). This classical view has, however, overlooked new attack surfaces that arise in more recent FL paradigms that instead communicate and aggregate external adaptation modules of pre-trained foundation models, e.g. soft-prompts prepended to input embeddings. (Weng et al., 2024).
To raise awareness of the potential existence of unexplored privacy threats in recent FL paradigms, this paper reveals a novel attack surface via manipulating the clustering and aggregation of local sets of soft prompts in federated prompt-tuning (FPT) (Weng et al., 2024). To the best of our knowledge, this is the first work that identifies an unexplored privacy risk in soft prompt optimization for representation fine-tuning, particularly in FPT. This is orthogonal to and complement existing research on privacy risks in prompting large language models with exemplars and/or task instructions that contain private data for in-context learning (ICL).
Such studies focus largely on designing adversarial queries that expose whether a particular data point was included in private instruction prompts used for in-context learning (Wen et al., 2024; Duan et al., 2023). In contrast, soft prompt optimization does not include private data in the prompts themselves. Moreover, the attack model in in-context learning originates from end users external to the training system whereas in federated prompt-tuning, it arises from the server within the training network. Existing defenses against prompt risk in in-context learning (Wu et al., 2023; Hong et al., 2023; Tang et al., 2023) are therefore not applicable to soft prompt risk in federated prompt-tuning.
To highlight the privacy risk that arises within soft prompt aggregation in FPT, we develop PromptMIA, a novel membership inference attack (MIA) that exploits FPT’s prompt update and selection mechanism as a new attack surface. This is substantiated with the following contributions:
(I) We develop an algorithm that optimizes adversarial prompts to be preferentially selected and updated by local clients when their datasets contain a target data point, while remaining unselected otherwise. PromptMIA is substantiated via running this algorithm on the server to create and add adversarial prompts to the shared prompt pools. The server can keep track of changes made to these prompts between communication iterations to infer membership information in a single communication round (Section 3.2.2).
(II) We analyze the theoretical attack advantage of PromptMIA from the lenses of a security game. In this view, the advantage of the attacker is characterized in terms of the true and false positive rates which measure the chance that the adversarial prompts are correctly and incorrectly selected, respectively. Our analysis shows that  PromptMIA achieves perfect true positive rate and provably low false positive rate, resulting in a high attack advantage (Section 3.2.3).
(III) We run extensive experiments to validate the effectiveness of PromptMIA across $7$ datasets including CIFAR-10, CIFAR-100, TinyImageNet, MNIST-M, Fashion-MNIST, CINIC-10, MMAFEDB. We also run validation experiments across $3$ vision transformer architectures including ViT-B/32 (Dosovitskiy et al., 2020), DeiT-B/16 (Touvron et al., 2021), and ConViT (d’Ascoli et al., 2021). Our empirical results consistently show that PromptMIA achieves $90\%$ attack success rate, which corroborates the aforementioned high attack advantage (Section 4.1).
(IV) We conduct additional analyses and experiments to demonstrate the effectiveness of PromptMIA against FL clients implementing standard defenses such as outlier detection, input noise perturbation, gradient obfuscation (Sections 4.2 and 4.3) and defenses that modify the prompt selection or aggregation protocol ( Appendix  A.4). We also note that PromptMIA exploits the prompt selection mechanism via monitoring which prompts are selected and updated rather than the content of the update, thus bypassing gradient obfuscation defenses (Duan et al., 2023). These results underscore the urgent need for more research on privacy defenses against this new prompt-based attack surface.
For clarity, we review the FPT framework in Section 2.
2 Federated Prompt-Tuning
In this section, we explain prompt-based learning and its extension to FPT. Prompt-based learning (Lester et al., 2021) reformulates the downstream task adaptation problem as input modification rather than weight updates. Given an image $x\in\mathbb{R}^{H\times W\times C}$ and a pretrained ViT model with frozen embedding layer $f_{e}$, let $x_{e}=f_{e}(x)\in\mathbb{R}^{L_{x}\times D}$. $L_{x}$ is the number of patches, and $D$ is the patch embedding’s dimension. A learnable prompt $P_{e}\in\mathbb{R}^{L_{p}\times D}$ is prepended to $x_{e}$ to form $x_{p}=[P_{e};x_{e}]$. A frozen attention stack $f_{a}$ followed by classification head $f_{c}$ produces predictions $\hat{y}=f_{c}\!\big(f_{a}(x_{p})\big).$ Learning to Prompt (L2P) (Wang et al., 2022) maintains a prompt pool of size $M$, denoted as $\mathcal{P}=\{P_{i}\in\mathbb{R}^{L_{p}\times D}\}_{i=1}^{M}$ with corresponding learnable keys $\mathcal{K}=\{k_{i}\in\mathbb{R}^{D_{k}}\}_{i=1}^{M}$. To facilitate query-key matching, a deterministic query function $q:\mathbb{R}^{H\times W\times C}\to\mathbb{R}^{D_{k}}$ is used. We use the pretrained model as a feature extractor and define the query feature as the [CLS] representation: $q(x)=f(x)[0,:].$ Using the cosine distance function $\gamma$, the top-$N$ prompts are chosen using Eq. 1:
$\hat{\mathcal{K}}_{x}$ $=\operatorname*{argmin}_{\{s_{i}\}_{i=1}^{N}\subseteq[M]}\sum_{i=1}^{N}\gamma\!\big(q(x),k_{s_{i}}\big)$ (1)
i.e., the $N$ prompts whose associated keys are closest to the query feature $q(x)$ under the cosine distance $\gamma$, or equivaliently, the $N$ prompts whose associated keys have the highest cosine similarity to $q(x)$. The adapted input is $x_{p}=[P_{s_{1}};\cdots;P_{s_{N}};x_{e}].$ Let the average of the prompt-token hidden states be $\bar{h}(x_{p})$ and the trainable classifier ${f_{c}}^{\phi}$ be parameterized by $\phi$, the training objective is defined as:
$\min_{\mathcal{P},\mathcal{K},\phi}\;\mathcal{L}\!\big(f_{c}^{\phi}(\bar{h}(x_{p})),y\big)+\lambda\!\sum_{i\in\hat{\mathcal{K}}_{x}}\gamma\!\big(q(x),k_{i}\big).$ (2)
The first term is the softmax cross-entropy, while the second is a surrogate that pulls the selected keys to align with the corresponding query features. The server keeps a global prompt pool $\mathcal{P}_{\textsc{global}}=\{P_{i}\}_{i=1}^{M}$. Given input image $x$, the client selects $\hat{\mathcal{K}}_{x}$ using Eq. 1 and updates $\hat{\mathcal{K}}_{x}$, the associated prompts $\hat{\mathcal{P}}_{x}$ along with ${f_{c}}^{\phi}$ according to Eq. 2. The server aggregates the trained prompts from participating clients to update global key-prompt pool $(\mathcal{K}_{\textsc{global}},\mathcal{P}_{\textsc{global}})$ using any of the existing FL algorithms (e.g., FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020)) or PFPT (Weng et al., 2024). See Appendix A.1 for more details.
Figure 1: PromptMIA workflow: (1) the server injects adversarial prompts designed for a target sample into the global prompt pool; (2) the modified pool is broadcast to clients; (3) each client performs query-key matching, which selects all adversarial prompts if the target sample is present; (4) selected prompts are locally updated and (5) returned to the server. By monitoring which prompts are updated, the server infers the target’s membership in client data.
3 MIAs against Federated Prompt-Tuning
In this section, we present the formulation and workflow of PromptMIA. An overview of the attack is shown in Fig. 1. Unlike prior approaches that rely on gradients, model weights, or auxiliary datasets to train shadow models, PromptMIA operates by injecting adversarial prompts $\mathcal{P_{\textsc{adv}}}$ into the shared prompt pool $\mathcal{P_{\textsc{global}}}$ prior to a training round. These prompts are associated with adversarial key vectors $\mathcal{K_{\textsc{adv}}}$ and are designed to be selected only when target sample $\mathcal{T}$ is present in a client’s data. As clients select and update prompts (Eq. 1), the server can monitor which injected prompts are modified and use this as a membership signal. This method exploits prompt updates as a privacy channel to infer membership in a single round, reducing both computational cost and adversarial assumptions.
3.1 Prompt based AMIs as Security Games
[…]
${\mathbf{\bf Adv}}^{\textup{AMI}}(\mathcal{A})=\Pr[b^{\prime}=1|b=1]-\Pr[b^{\prime}=1|b=0]$ (3)
where $b$ is the true membership label, and $b^{\prime}$ is the adversary’s prediction. We define the attack success rate as:
$\textbf{ASR}^{\textup{AMI}}(\mathcal{A})$ $=\tfrac{1}{2}\bigl(1+{\mathbf{\bf Adv}}^{\textup{AMI}}(\mathcal{A})\bigr)$ (4)
$=\tfrac{1}{2}\bigl(\Pr[b^{\prime}=1\mid b=1]+\Pr[b^{\prime}=0\mid b=0]\bigr)$
(a) t-SNE of global keys without adversarial key injection.
(b) t-SNE visualization shows naively injected adversarial keys overlapping
(c) t-SNE of global keys after PromptMIA adversarial keys injection.
Figure 2: Comparison of the global key distributions produced by a ViT-B/32 model trained on CIFAR-10 after 60 global epochs, visualized using t-SNE. Blue keys are benign keys, and red keys are adversarial keys.
3.2 MIA using Adversarial Prompt Injection
Given the security game laid out in Section 3.1, the adversary’s objective is to inject a set of $N$ adversarial keys-prompts $(\mathcal{K}_{\textsc{adv}},\mathcal{P}_{\textsc{adv}})$ into the global prompt pool such that, if $\mathcal{T}\in\mathcal{D}$, the top-$N$ selected prompts will always be the adversarial set. The membership signal is defined as the event that all adversarial prompts are selected and subsequently updated. To build intuition, we first introduce a Naive Prompt Injection attack and analyze its weaknesses, before introducing the more robust method, PromptMIA.
3.2.1 Naive Prompt Injection Attack
In this approach, the adversary constructs adversarial keys $\mathcal{K}_{\textsc{adv}}=\{\,k_{a_{m}}\,\}_{m=1}^{N}$ and inserts them into the global prompt pool. The goal is to ensure that if target data $\mathcal{T}\in\mathcal{D}$, the top-$N$ selected prompts using Eq. 1 always coincide with $\mathcal{K}_{\textsc{adv}}$. A straightforward strategy is to align each adversarial key $k_{a_{m}}$ directly with the client’s query vector $q(\mathcal{T})$ so as to maximize cosine similarity. We ensure that:
$\kappa\!\left(q(\mathcal{T}),k_{a_{m}}\right)=1\;\text{or}\;k_{a_{m}}=q(\mathcal{T}),\quad\forall\,k_{a_{m}}\in\mathcal{K}_{\textsc{adv}}.$ (5)
where $\kappa$ is the cosine similarity operator. The server selects a subset $S\subseteq[M]$ with $|S|=N$, and for each $j\in S$ performs key modificaton:
$(k_{j},P_{j})\;\mapsto\;(k_{a_{m}},P_{j}),\qquad m=1,\dots,N,$ (6)
However, this naive attack suffers from fundamental weaknesses. Since all adversarial keys collapse to the same vector $q(\mathcal{T})$ due to Eq. 5, clients can easily detect this redundancy through dimensionality reduction techniques (e.g., t-SNE, PCA) (see Fig. 2(b)), which would reveal a tight cluster of identical keys, or simply by directly inspecting key values. In addition, simple defenses such as discarding any key whose similarity exceeds a suspicious threshold (e.g., excluding $\kappa(q(\mathcal{T}),k_{a_{m}})>1-r$ for small $r>0$) can render the attack ineffective. Finally, because all adversarial keys are identical, the attack suffers from a high false positive rate ( see Section 4.1) : even when the target sample is absent, adversarial prompts may still be selected whenever the similarity of one adversarial key ( and as a result, all adversarial keys) to some non-target query $q(x)$, $x\in\mathcal{D},\,x\neq\mathcal{T}$, is higher than all benign keys, or $\kappa\!\left(q(x),k_{a_{m}}\right)>\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa\!\left(q(x),k_{b}\right)$.
Figure 3: Left: t-SNE projection of the global keys and query vectors from the train set. Center: K-Means clustering of keys with queries. Right: Each cluster modeled as a spherical Gaussian distribution centered at a key. All visualizations are generated from a ViT-B32 model trained on CIFAR-10 for 60 global epochs.
3.2.2 PromptMIA
To build up on and overcome the limitations of the naive attack, we impose two requirements on adversarial key generation. First, adversarial keys $\mathcal{K}_{\textsc{adv}}$ must achieve higher cosine similarity with the target query $q(\mathcal{T})$ than any benign key $\mathcal{K}_{\textsc{benign}}$. Second, adversarial keys must be sufficiently diverse from one another to not be trivially exposed. Formally, the server selects a subset $S\subseteq[M]$ of size $N$, and for each index $j\in S$, the original key $k_{j}$ is replaced with an adversarial key $k_{a_{m}}\in\mathcal{K}_{\textsc{adv}}$ ( see Eq. 6). $k_{a_{m}}$ is constrained to lie within a cosine similarity interval with $q(\mathcal{T})$:
$\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa\!\left(q(\mathcal{T}),k_{b}\right)+\delta_{\min}\;\leq\;\kappa(q(T),k_{a_{m}})\;\leq$ (7)
$\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa\!\left(q(\mathcal{T}),k_{b}\right)+\delta_{\min}+\Delta,\quad\forall k_{a_{m}}\in\mathcal{K}_{\mathrm{adv}}.$
where $\delta_{\min}$ ensures that all adversarial keys have higher cosine similarity to $q(\mathcal{T})$ than benign keys, while $\Delta$ introduces controlled variability to prevent all adversarial keys from collapsing to the same vector. Algorithm 1 describes the procedure for generating a single adversarial key with a specified cosine similarity $s$ to the target query. Building on this primitive, Algorithm 2 constructs a set of $N$ adversarial keys $\mathcal{K}_{\textsc{ADV}}$ that satisfies Eq. 7. Details of these algorithms are given in Appendix A.2. PromptMIA allows the adversary to generate an adversarial key set $\mathcal{K}_{\textsc{adv}}$ that will always be selected and updated when $\mathcal{T}\in\mathcal{D}$, while not being easily detectable (see Fig. 2(c) and Section 4.2). The server generates and distributes the modified global prompt pool $(\tilde{\mathcal{K}},\tilde{\mathcal{P}})=\bigl(\mathcal{K}_{\textsc{adv}}\cup\mathcal{K}_{\textsc{benign}},\;\mathcal{P}_{\textsc{adv}}\cup\mathcal{P}_{\textsc{benign}}\bigr)$ to all participating clients. Given client $t$ with local dataset $\mathcal{D}$ and input data $x\in\mathcal{D}$, the client computes $q(x)$ and selects top-$N$ keys $\hat{\mathcal{K}}_{x}$ using Eq. 1 and update only the selected prompts using Eq. 2. If $x\equiv\mathcal{T}$, then we have the top-$N$ keys are exactly the adversarial keys $\mathcal{K}_{\textsc{adv}}$, or $\hat{\mathcal{K}}_{x}\equiv\mathcal{K}_{\textsc{adv}}$ , which corresponds precisely to the membership condition.
3.2.3 Theoretical Analysis of PromptMIA
In this section, we theoretically analyze the performance of PromptMIA. Specifically, we study the True Positive Rate (TPR), the adversary’s success in identifying the target sample $\mathcal{T}$ when it is a member ($b=1$), and the False Positive Rate (FPR) which is the adversary’s error in identifying $\mathcal{T}$ as a member when it is a non-member ($b=0$).
First, the PromptMIA attack is constructed to ensure perfect identification when the target sample is present. Theorem 3.1 establishes that the TPR of PromptMIA equals 1. The proof follows directly from the construction of adversarial keys in Algorithm 2, which ensures that $\min_{k_{a}\in\mathcal{K}_{\textsc{adv}}}\kappa(q(\mathcal{T}),k_{a})>\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa(q(\mathcal{T}),k_{b})$. Thus, if the client possesses $\mathcal{T}\in\mathcal{D}$, the top-$N$ selected keys must be exactly the set $\mathcal{K}_{\textsc{adv}}$.
Theorem 3.1 (True Positive Rate).
Let $\mathcal{K}_{\textsc{adv}}=\{k_{a_{m}}\}_{m=1}^{N}$ be the set of $N$ adversarial keys generated by Algorithm 2 with parameters $\delta_{\min}>0$ and $\Delta\geq 0$. Let $\mathcal{K}_{\textsc{benign}}$ be the set of MNM-N benign keys. If the client’s dataset $\mathcal{D}$ contains the target sample $\mathcal{T}$ (i.e., $b=1$) and the client’s selection mechanism (Eq. 1) selects the top-$N$ prompts based on highest cosine similarity (lowest cosine distance), the set of selected keys $\hat{\mathcal{K}}_{\mathcal{T}}$ for the query $q(\mathcal{T})$ will be exactly the adversarial set: $\hat{\mathcal{K}}_{\mathcal{T}}=\mathcal{K}_{\textsc{adv}}$. Consequently, the True Positive Rate (TPR) of PromptMIA is 1.
$TPR=\Pr[b^{\prime}=1\mid b=1]=1$
We now analyze the FPR, $\Pr[b^{\prime}=1\mid b=0]$, which is more complex. A false positive arises whenever a non-member query $q(x)$ selects all $N$ adversarial keys $\{k_{a_{j}}\}_{j=1}^{N}$ as its top-N choices than to any benign key $\{k_{b_{i}}\}_{i=1}^{M-N}$. To bound this probability, we first model the distribution of non-member query $q(x)$ relative to the benign keys. The training objective in Eq. 2 includes a surrogate loss $\gamma(q(x),k_{i})$ that explicitly pulls a selected key $k_{i}$ to align with the query feature $q(x)$ that selected it. Thus, after several rounds of training, the benign keys $\mathcal{K}_{\textsc{benign}}$ will stabilize and function as the centroids for the query vectors $q(x)$ generated from the clients’ data (see Fig. 3 and Fig. 15). This observation forms our foundational assumptions:
Assumption 1 (Benign Keys as Cluster Centroids). We assume that the K=MNK=M-N benign keys $k_{b_{i}}\in\mathcal{K}_{\textsc{benign}}$ act as the effective centroids of the non-member query vector distribution. Each non-member query $q(x)$ is assumed to belong to the cluster of its nearest benign key.
Following Assumption 1, we model the distribution of non-member queries $q(x)$ belonging to benign cluster $i$ in Assumption 2. This is a reasonable assumption since the prompt learning mechanism in the federated prompt-tuning framework models the prompt set as a sample from a Poisson point process with a Gaussian base measure and likelihood (Weng et al., 2024). The aggregated prompts serve as centroids of prompt clusters, which are optimized to lie close (on average) to different regimes of input queries in Euclidean space. Consequently, it is natural to expect the input queries to be partitioned into Gaussian clusters centered around these aggregated prompts. This clustering behavior has also been verified and visualized in Fig. 3.
Assumption 2 (Gaussian Query Model). We model the distribution of non-member queries $q(x)$ belonging to benign cluster $i$ as a spherical Gaussian distribution centered at that key, i.e., $q(x)\sim\mathcal{N}(k_{b_{i}},\sigma_{i}^{2}I)$ where $\sigma_{i}^{2}$ is the variance of the non-member queries associated with that key.
Next, we introduce the adversarial key into this setting. For a non-member query $q(x)$ drawn from cluster $i$, we define a cluster-flip event $E_{i}$ as the case when $q(x)$ selects all $N$ adversarial keys as its $N$ nearest centroids. Formally, $E_{i}$ is the intersection of $N\times(M-N)$ race events $A_{jl}$, where each $A_{jl}$ denotes that $q(x)$ is closer to an adversarial key $k_{a_{j}}$ than to a benign key $k_{b_{l}}$. A cluster-flip event thus arises only if the adversary wins all of these races simultaneously. The probability of the joint event $E_{i}$ can then be upper bounded by the probability of the single race with the lowest success probability, as formally stated in Lemma 3.2.
Lemma 3.2.
Let $q(x)\sim\mathcal{N}(k_{b_{i}},\sigma_{i}^{2}I)$ be a non-member query from benign cluster $i$. The probability $\Pr(E_{i})$, that $q(x)$ selects all $N$ adversarial keys $\mathcal{K}_{ADV}=\{k_{a_{1}},\dots,k_{a_{N}}\}$ as its $N$ closest centroids, is bounded by:
$\Pr(E_{i})\leq\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(\frac{(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}}{\sigma_{i}\|k_{a_{j}}-k_{b_{l}}\|}\right)$
where $\Phi(\cdot)$ is the CDF of the standard normal distribution.
With this lemma, we can bound the FPR. The FPR is the probability of the event $E_{FP}$ that that a single, random non-member query $q(x)$ results in a cluster-flip event. Using the Law of Total Probability, FPR can be expressed as a weighted sum of cluster-flip probabilities, where the weights correspond to the prior probabilities of clusters. Based on the bound on the probability of cluster-flip events for each cluster established in Lemma 3.2, we can derive a bound on the FPR which is the largest (worst-case) cluster-flip probability across all clusters, stated in Theorem 3.3.
Theorem 3.3 (False Positive Rate).
The per-sample False Positive Rate (FPR) is bounded by:
$FPR\leq\max_{1\leq i\leq M-N}\left(\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(z_{ijl}\right)\right)$
where $\mathcal{K}_{ADV}=\{k_{a_{1}},\dots,k_{a_{N}}\}$ is the set of $N$ adversarial keys, and $z_{ijl}$ is the $z$-score: $z_{ijl}=\frac{(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}}{\sigma_{i}\|k_{a_{j}}-k_{b_{l}}\|}$
Theorem 3.3 provides some insights on conditions under which the attack is most effective. First, the bound is tighter when the adversarial target $q(\mathcal{T})$ is highly distinctive. A distinctive target ensures that its corresponding adversarial keys are geometrically well separated from benign key clusters, resulting in small $\Phi\left(z_{ijl}\right)$. In addition, the FPR is lower when the benign data forms tight and compact clusters in the query space around the benign keys (i.e., minimizing $\sigma_{i}^{2}$). That will decreases the likelihood that any non-member query will randomly stray into the adversary’s region. Additionally, combining Theorems 3.1 and 3.3 yields the bound on the Advantage stated in Corollary 3.4. The detailed proofs are provided in Appendix A.7.
Corollary 3.4 (Attack Advantage).
The advantage of the adversarial server $\mathcal{A}$, which is defined as ${\mathbf{\bf Adv}}^{\textup{AMI}}(\mathcal{A})=TPR-FPR$, is lower bounded by:
${\mathbf{\bf Adv}}^{\textup{AMI}}(\mathcal{A})\geq 1-\max_{1\leq i\leq M-N}\left(\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(z_{ijl}\right)\right)$
Figure 4: Performance of PromptMIA vs Naive averaged across three models. Each subplot shows Advantage and Attack Success Rate w.r.t Batch Size across CIFAR10, CIFAR100, TinyImageNet, and FourDataset.
(a) Isolation Forest
(b) Local Outlier Factor
(c) OneClassSVM
(d) EllipticEnvelope
Figure 5: Visualization of outlier detection methods on CIFAR-10 trained Vit-B32. Blue keys are benign keys. Red keys are adversarial keys. Crossed keys are flagged as outliers from the corresponding algorithm.
3.3 MIA Defenses against PromptMIA
Although the goal of this work is not to develop new defense mechanisms, we examine how standard approaches originally designed for gradient or output based MIAs interact with PromptMIA. A widely used defense is Differentially Private SGD (DPSGD) (Abadi et al., 2016; Duan et al., 2023), which clips per-sample gradients and injects noise before aggregation to provide $(\varepsilon,\delta)$-privacy guarantees. While DPSGD protects the content of gradients, in federated prompt tuning the top-$N$ prompt selection is independent of the gradient update procedure and therefore remains exposed, rendering DPSGD ineffective against PromptMIA. Therefore, we do not further evaluate DPSGD in our experiments. We consider Input Noise Perturbation, which adds calibrated noise directly to input pixels (Lecuyer et al., 2019). Similar to DPSGD, this approach incurs privacy-utility trade-off. Another line of defense is to use Anomaly Detection Methods (IsolationForest (Liu et al., 2008), LocalOutlierFactor (Breunig et al., 2000), OneClassSVM (Manevitz and Yousef, 2001), and EllipticEnvelope (Rosseeuw, 1999) to filter out adversarial prompts. The effectiveness of anomaly detection and input noise perturbation against PromptMIA are reported in Sections 4.2 and 4.3, respectively. Although it would be interesting to evaluate deep learning-based anomaly detection algorithms (Do et al., 2025; Jiang et al., 2023) against PromptMIA, we leave this for future work. Moreover, the number of prompts in the prompt pool is typically small, which raises doubts about the effectiveness of deep learning based methods in this setting. Appendix A.3 presents a detailed discussion on theses defenses. We also analyzes the implications of other potential system-level defense strategies such as randomized key indices, secure aggregation (Bonawitz et al., 2017), and randomized prompt selection via prompt dropout and show that these defenses are either provably or empirically ineffective (Appx. A.4).
While our results (Section 4.2) show that traditional outlier detection fails to reliably identify adversarial keys, we further introduce a hyperparameter $\beta$ to control the alignment between adversarial and benign keys. Larger $\beta$ improves PromptMIA robustness against stronger anomaly detectors, at the cost of reduced MIA accuracy (Appendix A.5).
4 Experimental Results
We evaluate PromptMIA on four datasets: CIFAR-10, CIFAR-100  (Krizhevsky et al., 2009), TinyImageNet  (Le and Yang, 2015), and a synthetic 4-dataset benchmark constructed by pooling MNIST-M (Lee et al., 2021), Fashion-MNIST  (Xiao et al., 2017), CINIC-10  (Darlow et al., 2018), and MMAFEDB  [^1] https://www.kaggle.com/datasets/yuulind/mmafedb-clean and three different models: ViT-B32, ConViT and DeiT. Experiments follow four axes: (1) measuring attack effectiveness via advantage and success rate; (2) testing robustness of PromptMIA against classical anomaly detection methods; (3) analyzing the impact of input noise perturbation defenses; and (4) conducting ablations on key hyperparameters ($M$, NN, $\beta$, $\delta_{\min}$, $\Delta$). Details about experimental settings are given in Appx. A.6.
4.1 Advantage and Attack Success Rate Measurement
We evaluate the performance of PromptMIA against Federated Prompt Tuning using two metrics: Advantage (Eq. 3) and Attack Success Rate (Eq. 4). For all experiments, we set $\delta_{\text{min}}=0.02$ and $\Delta=0.05$. Unless otherwise specifically noted, we set $\beta=0$. Following (Wang et al., 2022), the global prompt pool size is fixed at $M=20$, and the prompt selection size at $N=4$. In the case of batched update, given batch $B=\{(x_{i},y_{i})\}_{i=1}^{\ell}$, for each sample the client computes the per-sample selected key set $\hat{\mathcal{K}}_{x_{i}}$ and corresponding per-sample loss $\mathcal{L}_{x_{i}}$ (defined in Eq. 2) and updates the batch-level set of chosen keys and prompts: $\hat{\mathcal{K}}_{B}=\hat{\mathcal{K}}_{B}\bigcup\hat{\mathcal{K}}_{x_{i}}$ and $\hat{\mathcal{P}}_{B}=\hat{\mathcal{P}}_{B}\bigcup\mathcal{P}_{x_{i}}$. Batch-wise loss $\mathcal{L}_{B}$ is calculated by accumulating $\mathcal{L}_{x_{i}}$. The client update the selected keys and prompts as $\hat{\mathcal{K}}_{B}\leftarrow\mathcal{K}_{B}-\mu\nabla_{\hat{\mathcal{K}}_{B}}\mathcal{L}_{B}$ and $\hat{\mathcal{P}}_{B}\leftarrow\hat{\mathcal{P}}_{B}-\mu\nabla_{\hat{\mathcal{P}}_{B}}\mathcal{L}_{B}$, where $\mu$ is the learning rate. After receiving the client’s updates, the server infers membership by checking whether all adversarial prompts are selected, i.e. $\mathbb{I}_{\{\mathcal{T}\in\mathcal{D}\}}=1$ if $\mathcal{K}_{\textsc{ADV}}\subseteq\hat{\mathcal{K}}_{B}$, and 0 otherwise. Figure 4 shows average results of three models over four different datasets. Detailed results on individual models are given in Appendix A.8. PromptMIA consistently achieves near-perfect attack success rates across all models and datasets at small batch sizes, and maintains $90\%$ ASR and Advantage at larger batch sizes. In contrast, naive prompt injection collapses as the batch size increases, reaching Advantage of only $\approx 20\%$ against FourDataset when batch size =256=256.
4.2 Performance of Outlier Detection Dethods
To evaluate outlier detection methods against PromptMIA, we frame adversarial key detection as an unsupervised anomaly-detection task over the global prompt pool. For each input $x\in\mathcal{D}$, we flip $\tilde{b}\sim\mathrm{Bernoulli}(1/2)$: if $\tilde{b}=1$, $N$ adversarial keys are injected into the pool $\tilde{\mathcal{K}}$; otherwise, the pool remains unmodified (clean control). We then apply IsolationForest, LocalOutlierFactor, OneClassSVM and EllipticEnvelope to score each key, where keys identified as outliers are considered adversarial. Results are averaged across all datasets and models (detailed results in App. A.9).
(a) CIFAR10
(b) CIFAR100
(c) TinyImagenet
(d) FourDataset
Figure 6: Attack Success Rate of PromptMIA under Input Noise Perturbation with different $\epsilon$.
Table 1: Outlier detection results averaged over all datasets and models.
Method Precision Recall F1
IsolationForest 0.2672 1.0000 0.4172
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.2038 0.4993 0.2851
EllipticEnvelope 0.0024 0.0052 0.0033
Table 2: Model accuracy (%) under local differential privacy with different privacy budgets $\epsilon$ using ViT-B32.
Dataset $\epsilon=3$ $\epsilon=5$ $\epsilon=8$ Non-DP
CIFAR-10 0.85 0.88 0.90 0.95
CIFAR-100 0.50 0.59 0.62 0.78
TinyImageNet 0.72 0.76 0.79 0.86
FourDataset 0.55 0.59 0.64 0.76
Table 1 shows that naively applying anomaly detection is ineffective against PromptMIA. While IsolationForest achieves high recall, it does so by broadly flagging almost all benign keys as adversarial. LocalOutlierFactor and EllipticEnvelope detected almost none of the injected adversarial keys. OneClassSVM achieves moderate recall but also suffers from a high false positive rate. These findings highlight the ineffectiveness of traditional outlier detection methods against PromptMIA. Figures 5 and 13 provide visualization of outlier detection methods against PromptMIA.
4.3 Performance and Impact of Noise Perturbation
We evaluate the effectiveness of input noise perturbation as a defense against PromptMIA. In particular, we measure the attack success rate (ASR) of PromptMIA under three privacy budgets, $\epsilon\in\{3,5,8\}$, using the ViT-B32 model across four different datasets, as shown in Fig. 6. Table 2 reports the privacy-utility trade-off. We find that using a larger $\delta_{\min}$ ($0.2$) works better under input noise perturbation. If $\delta_{\min}$ is too small, the adversarial keys are only slightly closer to $q(\mathcal{T})$ than the benign keys, so even a small amount of noise can break the attack. As the batch size grows, increasing noise begins to reduce the attack’s success, but only when very strong privacy guarantees are applied. For instance, on CIFAR-100 with $\epsilon=3$, the ASR decreases noticeably at larger batch sizes, but this protection comes at a substantial cost in accuracy (dropping from $0.78$ to $0.50$). At moderate privacy levels ($\epsilon=5$ and $\epsilon=8$), input noise fails to offer meaningful protection: the attack continues to achieve moderate to high ASR on three out of four datasets, even when the batch size is large. These results highlight the trade-off: strong noise can partially suppress PromptMIA but severely harms accuracy, while weaker noise preserves accuracy but leaves the model more vulnerable.
4.4 Ablation Experiments
We analyze how key hyperparameters (global pool size $M$, selection size $N$, $\delta_{\min}$, $\Delta$, $\beta$, and number of training rounds) affect PromptMIA’s performance. The most interesting finding is that the attack success rate of PromptMIA is much higher on the models that have been trained for a few epochs rather than randomly initialized keys (Fig. 14(f)). This corresponds to our theoretical findings in Section 3.2.3 that FPR is lower when the benign data forms tight and compact clusters in the query space around the benign keys, which happens naturally during the training process ( Appx. A.11). More detailed analysis and insights on hyperparameters are provided in Appendix A.10. We provide additional experiments under very large batch size in Appendix  A.12. To show that our attack generalizes to to all variants of FPT that adopt the common paradigm of a frozen backbone model (often transformer-based) paired with a shared, learnable (soft) prompt pool across clients, we provide additional experiments on multimodal and text data in Appendix  A.13. Additional studies on the attack success rate and distribution of global keys and benign query vectors under heterogeneous settings is given in Appendix  A.14.
5 Related Work
Membership Inference Attacks against Federated Learning. The goal of MIAs in FL is to identify if a specific data point was part of a client’s training set. Passive attacks (Shokri et al., 2017) involve an honest-but-curious server observing the model updates, while Active Membership Inference (AMI) attacks involve a dishonest server poisoning the global models, e.g., maliciously modifying model parameters, before dispatching them to clients (Nguyen et al., 2023; Vu et al., 2024). Recently, (Zhu et al., 2025) proposed a three-step attack that leverages updates from all clients that can be integrated as an extension to existing attacks.
Federated Fine-Tuning of Foundation Models. Federated fine-tuning avoids updating the entire model and instead allows clients to fine-tune only a small subset of parameters. These include LoRA-based methods (Qi et al., 2024; Wang et al., 2024; Fan et al., 2025), adapter-based approaches (Cai et al., 2022; Ghiasvand et al., 2024). Recently, prompt-based federated fine-tuning approaches have been proposed (Su et al., 2024; Weng et al., 2024; Bai et al., 2024; Feng et al., 2024), which update soft prompts instead of full model weights during FL. While it is possible that LoRA and adapter-based federated fine-tuning may also exhibit similar vulnerabilities, our work focuses on MIAs under FPT setting since these approaches are architecturally orthogonal.
Privacy Risks in LLM Prompting. LLMs are strong in-context learners that can adapt to downstream tasks by prepending discrete prompts such as exemplars or task instructions without requiring fine-tuning. These exemplars often contain sensitive information (e.g., medical records, personally identifiable data). Adversaries can exploit this by crafting malicious prompts to extract confidential information in these discrete prompts (Wen et al., 2024; Duan et al., 2023). To mitigate such risks, recent work has proposed a range of defense strategies primarily based on the notion of differential privacy (Duan et al., 2023; Wu et al., 2023; Hong et al., 2023; Tang et al., 2023). Our work is the first to investigate the privacy risks of soft prompts in FPT.
6 Conclusion
In conclusion, we show that FPT introduces a new and critical privacy vulnerability. Through PromptMIA, we demonstrate that a malicious server can leverage adversarial prompts to reliably infer client membership, achieving high attack success rates across multiple datasets. Our theoretical analysis explains the attack’s robustness by establishing a lower bound on the attack’s advantage. Finally, evaluation of PromptMIA against existing MIA defenses reveals their limitations in this setting, underscoring the need for new defense strategies specifically designed for FPT.
Impact Statement
This work reveals that federated prompt tuning introduces new privacy risks, showing that active membership inference attacks can compromise client data under practical federated learning settings. Our results demonstrate that existing defenses are either ineffective or incur prohibitive utility costs to provide meaningful protection. These findings have important implications for deploying federated prompt-tuning in privacy-sensitive domains such as healthcare and finance, underscoring the need for prompt-aware privacy safeguards and more robust defenses. We hope this work motivates further research on secure and privacy-preserving prompt-based federated learning.
Appendix A Appendix
A.1 Federated Prompt Tuning
Figure 7 illustrates the workflow of federated prompt tuning. In federated prompt tuning, a central server maintains a global pool of prompts and keys. Each client selects the top-$N$ prompts most similar to its input features, updates those prompts and a lightweight classifier locally, and sends the updated prompts back. The server then aggregates the clients’ prompts to refine the global prompt pool without sharing raw data.
Figure 7: (Local Phase) each client samples and fine-tunes a subset of global summarizing prompts using a prompt-selection strategy; (Global Phase) the server aggregates all local prompt sets to refine the global prompt pool.
A.2 Generating adversarial keys set $\mathcal{K}_{\textsc{adv}}$
To mount PromptMIA, the adversarial server must generate a set of keys $\mathcal{K}_{\textsc{ADV}}$ that are more similar to the target query vector $q(\mathcal{T})$ than all benign keys $\mathcal{K}_{\textsc{benign}}$ while remaining diverse enough to avoid detection.
GenKeyWithSimilarity (Alg. 1) constructs a single adversarial key with a desired cosine similarity $s$ to the target query. It first normalizes the target query $q(\mathcal{T})$ to obtain the unit vector $\hat{q}=q(\mathcal{T})/\|q(\mathcal{T})\|$. Then it samples a random vector $r\in\mathbb{R}^{D_{k}}$ and removes the component of $r$ that lies along $\hat{q}$ by computing $o=r-\langle r,\hat{q}\rangle\hat{q}$. This ensures that $o$ is orthogonal to the target direction. Next, $o$ is normalized to $\hat{o}=o/\|o\|$, producing a unit vector exactly perpendicular to $\hat{q}$. To construct a vector with a desired cosine similarity $s$ to $\hat{q}$, the algorithm forms
$\hat{v}=s\cdot\hat{q}+\sqrt{1-s^{2}}\cdot\hat{o}$,
Finally, $\hat{v}$ is rescaled to match the original norm of $q(\mathcal{T})$, producing the adversarial key $k_{a}=\hat{v}\cdot\|q(\mathcal{T})\|$.
GenAdvKeySet (Alg. 2) builds the entire adversarial key set $\mathcal{K}_{\textsc{ADV}}$. It computes the maximum similarity $s_{\max}$ between the target query and existing benign keys, then samples $N$ similarity scores uniformly from the interval
$s_{\max}+\delta_{\min},\;s_{\max}+\delta_{\min}+\Delta],$
ensuring that each adversarial key is slightly closer to the target query than any benign key by at least $\delta_{\min}$ but not so close that all keys collapse to the target query. Each sampled similarity $s_{m}$ is used to generate a key via GenKeyWithSimilarity, producing a diverse set of keys that satisfy the required similarity bounds.
[…]
0: Target query vector $q(\mathcal{T})\in\mathbb{R}^{D_{k}}$, desired cosine similarity $s\in(0,1)$
0: Adversarial key $k_{a}\in\mathbb{R}^{D_{k}}$ such that $\kappa(k_{a},q(\mathcal{T}))\approx s$ and $\|k_{a}\|=\|q(\mathcal{T})\|$
[…]
7:return $k_{a}$ {Adversarial key}
[…]
0: Target query vector $q(\mathcal{T})\in\mathbb{R}^{D_{k}}$, benign key set $\mathcal{K}_{\textsc{benign}}$, margins $\delta_{\min},\Delta$, number of adversarial keys NN
[…]
3:for $m=1$ to NN do
4:  Sample $s_{m}\sim\mathcal{U}(s_{\max}+\delta_{\min},\;s_{\max}+\delta_{\min}+\Delta)$
[…]
6:end for
7:return Adversarial key set $\mathcal{K}_{\textsc{adv}}=\{k_{a_{m}}\}_{m=1}^{N}$
A.3 Membership Inference Defenses
While the focus of our work is not on designing new defenses, we discuss how standard approaches originally developed for gradient-based or output-based MIAs can be adapted to, and interact with, PromptMIA.
PromptDPSGD. A widely used defense is Differentially Private SGD (DPSGD), which clips per-sample gradients and injects noise before aggregation to provide $(\varepsilon,\delta)$-privacy guarantees. Such methods protect the content of gradients associated with updated prompts. However, in federated prompt tuning, each client still reveals which top-$N$ prompts it selects and updates, since this prompt selection mechanism is independent of the gradient update procedure. Thus, while DPSGD masks gradient values, it leaves selection patterns unchanged and thus this defense is ineffective against PromptMIA. DPSGD also incurs a significant privacy–utility trade-off  (Abadi et al., 2016).
Noise Perturbation Rather than obfuscating gradient updates, clients can achieve differential privacy (DP) w.r.t the input by injecting calibrated noise directly into the input image (Lecuyer et al., 2019). This introduces randomness to Eq. 1:
$\mathcal{K}_{x}=\operatorname*{argmax}_{\{s_{i}\}_{i=1}^{N}\subseteq[M]}\sum_{i=1}^{N}\kappa\!\big(q(x+\eta),k_{s_{i}}\big),\qquad\eta\sim\mathcal{N}(0,\tilde{\sigma}I).$ (8)
Increasing the noise variance strengthens the protection by making the prompt selection less predictable and reducing the effectiveness of adversarially crafted keys. However, this also comes at the high cost of model utility.
Anomaly detection techniques We test commonly used anomaly detection techniques in machine learning against PromptMIA. Specifically, we consider the following classical approaches: IsolationForest (Liu et al., 2008), LocalOutlierFactor (Breunig et al., 2000), OneClassSVM (Manevitz and Yousef, 2001), and EllipticEnvelope (Rosseeuw, 1999). These methods are widely used, well established in the literature, and directly available in scikit-learn library (Kramer, 2016). Although it would be interesting to evaluate deep learning-based anomaly detection algorithms (Do et al., 2025; Jiang et al., 2023) (e.g., autoencoders, VAEs, GAN-based models) against PromptMIA, we leave this for future work. Moreover, the number of prompts in the prompt pool is typically small (20 in our experiments), which raises doubts about the effectiveness of deep learning based methods in this setting. We therefore restrict our evaluation to the aforementioned classical methods. We briefly describe the anomaly detection algorithms used in this paper:
  • Isolation Forest detects anomalies by recursively partitioning the feature space with random splits. Each sample’s path length, or the number of splits needed to isolate it in a random tree is shorter for outliers, as they are easier to separate from the bulk of the data. Averaging this path length over a forest of random trees yields an anomaly score: points with shorter average paths are more likely to be anomalous.
  • Local Outlier Factor (LOF) detects anomalies by comparing the local density of each sample to that of its $k$-nearest neighbors. Normal points have similar density to their neighbors, while outliers lie in sparser regions. The LOF score is the ratio of the average neighbor density to the sample’s own density; values $\text{LOF}\!>\!1$ indicate potential outliers. LOF captures both local and global structure, making it effective in datasets with varying densities.
  • One-Class Support Vector Machine (OCSVM) is an unsupervised anomaly detection method derived from the Support Vector Machine framework. Instead of separating multiple classes, OCSVM learns a decision boundary that encloses the majority (normal) data in feature space, labeling points outside this region as anomalies or novelties. It works by maximizing the margin around normal data to create a robust “normalcy region” using a kernel function (commonly the radial basis function) to capture non-linear patterns.
  • Elliptic Envelope is an outlier detection method that assumes inliers follow a known distribution, typically Gaussian. It fits a robust estimate of the data’s mean and covariance (using the Minimum Covariance Determinant estimator) to capture the central elliptical shape of normal data while ignoring outliers. Points are then scored by their Mahalanobis distance from this fitted ellipse, with distant points flagged as anomalies. This approach is effective when the normal data distribution is approximately Gaussian.
Figure 8: Impact of Prompt Dropout on attack success rate.
A.4 System-level Defenses
We also analyze the impact of system-level defenses that modify the prompt selection or aggregation mechanism against PromptMIA.
  • Randomized key indices. We consider the setting where, after training, clients randomly permute the indices of their prompt keys before sending updates to the server to prevent the server from knowing exactly which prompts were updated.  However, this defense is fundamentally ineffective against PromptMIA. The server already stores the previous prompt pool and can check whether each adversarial prompt appears in the client-updated pool via content matching, which is unaffected by index permutation. It then knows that the client selected and updated all adversarial prompts when no match is found. Index randomization is thus provably ineffective as it cannot prevent the server from knowing which prompts were updated.
  • Secure Aggregation. Another strategy is to apply secure aggregation (Bonawitz et al., 2017) so that individual client updates are hidden from the server. The intuition is that, if the server only observes an aggregated result rather than each client’s prompt updates, it may be unable to determine whether a particular adversarial prompt was selected. However, current secure aggregation protocols are developed for linear aggregation (that is, weighted averaging) of local model parameters, and do not extend to probablistic aggregation methods commonly used in federated prompt tuning. Weighted averaging is not sufficient in non-IID settings where local parameters must first be aligned before aggregation in order to prevent important features from collapsing into less informative representations due to semantic misalignment (Weng et al., 2024). To address this, the recent PFPT work (Weng et al., 2024) developed an aligned and aggregated mechanism based on probabilistic non-parametric clustering which was shown to achieve substantially better performance than weighted averaging in non-IID settings. However, both the alignment and the aggregation steps in PFPT are inherently non-linear, making it unclear how existing secure aggregation methods could be generalized to support such a mechanism.
  • Prompt Dropout. Since PromptMIA relies on all adversarial keys being selected, introducing randomness into the prompt selection (e.g via prompt dropout) can reduce the attack success rate but it remains substantial (larger than 75%) as shown in our experiment with prompt dropout in Fig. 8.
A.5 Controlling the alignment of adversarial and benign keys
While our experimental results (Section 4.2) show that it is not trivial to use traditional anomaly detection algorithms to detect adversarial prompts generated by PromptMIA, we also propose and extension to improve the stealthiness of PromptMIA in case a stronger anomaly detection algorithm is used by introducing a hyperparameter $\beta$ that controls the alignment of $\mathcal{K_{\textsc{adv}}}$ and $\mathcal{K_{\textsc{benign}}}$. In particular, we make modification to Alg. 1 as follow:
[…]
0: Target query vector $q(\mathcal{T})\in\mathbb{R}^{D_{k}}$, desired cosine similarity $s\in(0,1)$, benign key set $\mathcal{K}_{\textsc{benign}}$, mixing factor $\beta\in(0,1)$
0: Vector $k_{a}\in\mathbb{R}^{D_{k}}$ such that $\kappa(k_{a},q(\mathcal{T}))\approx s$ and $\|k_{a}\|=\|q(\mathcal{T})\|$
[…]
3:Sample $k_{b}\sim\mathcal{K}_{\textsc{benign}}$ {Random benign key from the set}
[…]
11:return $k_{a}$
Alg. 3 extends Alg. 1 by introducing a mixing factor $\beta$ that interpolates between a random direction and a sampled benign key. By adjusting $\beta$, the adversarial key is made statistically closer to benign keys while still achieving the target cosine similarity $s$ with the query. Specifically, small values of $\beta$ result in keys that are more random and thus easier to separate from benign ones, whereas larger values of $\beta$ increase alignment with benign keys. This alignment improves the stealthiness of the attack, since anomaly detectors are more likely to misclassify adversarial keys as benign, but it also raises the false positive rate (FPR) of the attack because the top-$N$ selection mechanism may incorrectly select adversarial keys even when the target data $\mathcal{T}\notin\mathcal{D}$. Setting $\beta=0$ reduces Alg. 3 to Alg. 1. Experimental results on the attack success rate of PromptMIA under various $\beta$ is given in Section A.10. Figure 9 illustrates how $\beta$ can control the alignment between adversarial and benign keys.
(a) $\beta=0$
(b) $\beta=0.4$
(c) $\beta=0.6$
(d) $\beta=0.8$
Figure 9: t-SNE visualization of global keys after PromptMIA injection with different $\beta$ value. Adversarial keys are colored in red, while benign keys are colored in blue.
A.6 Experimental Settings
Datasets.
We evaluate our methods on four widely used vision benchmarks: CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009), TinyImageNet (Le and Yang, 2015), and a synthetic benchmark referred to as 4-dataset. The 4-dataset benchmark is constructed by pooling four diverse datasets: 1) MNIST-M (Lee et al., 2021), 2) Fashion-MNIST (Xiao et al., 2017), 3) CINIC-10 (Darlow et al., 2018), and 4) MMAFEDB [^2] https://www.kaggle.com/datasets/yuulind/mmafedb-clean. For 4-dataset, we use a total of 120,000 training and 10,000 test samples, with 30,000 training and 2,500 test examples drawn from each dataset, ensuring uniform class distributions. For CIFAR-10, CIFAR-100, and TinyImageNet, we adopt the standard train/test splits provided by the official datasets.
Federated learning setup.
To simulate heterogeneous client data, we partition datasets using a Dirichlet distribution with concentration parameter $\alpha=0.5$, which produces non-i.i.d. label distributions across clients. We consider a federation of 80 clients, with 10 randomly selected in each communication round. Local updates are performed with the Adam optimizer (learning rate $1\!\times\!10^{-4}$). Training is conducted for 60 communication rounds. We set hyperparameter $\lambda$ in eq. 2 to be 0.5, consistent with (Wang et al., 2022). We use three different baseline pretrain backbone: ViT-B/32 (Dosovitskiy et al., 2020), DeiT-B/16 (Touvron et al., 2021), and ConViT (d’Ascoli et al., 2021).
Evaluation protocol.
We organize our evaluation along three main dimensions. First, we measure the performance of PromptMIA in terms of advantage ( Eq. 3) and attack success rate (Eq. 4) in Section 4.1. Second, we assess the robustness of PromptMIA against classical anomaly detection methods such as Isolation Forest, Local Outlier Factor, One-Class SVM, and Elliptic Envelope (Section 4.2). Third, we study the effect of noise-based defenses on the input image on PromptMIA attack performance (Section 4.3). Finally, we perform ablation experiments to analyze the impact of current number of training round, data heterogeineity, input modalities and the role of the parameters $M$, NN, $\beta$, $\delta_{\min}$ and $\Delta$ on the performance of PromptMIA (Section 4.4).
Implementation details.
All experiments were conducted on a Linux workstation running Ubuntu 20.04 LTS, equipped with an Intel(R) Xeon(R) CPU E5-2697 v4 @ 2.30GHz (18 cores 36 threads), 384GB RAM, and two NVIDIA RTX A6000 GPU (48GB VRAM each). Our implementation is based on PyTorch 2.0 with CUDA 12.2.
A.7 Proofs
Theorem 3.1 (True Positive Rate).
Let $\mathcal{K}_{\textsc{adv}}=\{k_{a_{m}}\}_{m=1}^{N}$ be the set of $N$ adversarial keys generated by Algorithm 2 with parameters $\delta_{\min}>0$ and $\Delta\geq 0$. Let $\mathcal{K}_{\textsc{benign}}$ be the set of MNM-N benign keys. If the client’s dataset $\mathcal{D}$ contains the target sample $\mathcal{T}$ (i.e., $b=1$) and the client’s selection mechanism (Eq. 1) selects the top-$N$ prompts based on highest cosine similarity (lowest cosine distance), the set of selected keys $\hat{\mathcal{K}}_{\mathcal{T}}$ for the query $q(\mathcal{T})$ will be exactly the adversarial set: $\hat{\mathcal{K}}_{\mathcal{T}}=\mathcal{K}_{\textsc{adv}}$. Consequently, the True Positive Rate (TPR) of PromptMIA is 1.
$TPR=\Pr[b^{\prime}=1\mid b=1]=1$
Proof.
Based on Algorithm 2, every generated adversarial key $k_{a_{m}}\in\mathcal{K}_{\textsc{adv}}$ have a cosine similarity $s_{m}$ to $q(\mathcal{T})$ such that:
$s_{m}\geq s_{\max}+\delta_{\min},\qquad\text{where }s_{\max}=\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa(q(\mathcal{T}),k_{b})$
Since $\delta_{\min}>0$, we have:
$\min_{k_{a}\in\mathcal{K}_{\textsc{adv}}}\kappa(q(\mathcal{T}),k_{a})>\max_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\kappa(q(\mathcal{T}),k_{b})$
Because cosine distance $\gamma(\cdot,\cdot)$ is a monotonically decreasing function of cosine similarity $\kappa(\cdot,\cdot)$, we have an equivalent expression:
$\max_{k_{a}\in\mathcal{K}_{\textsc{adv}}}\gamma(q(\mathcal{T}),k_{a})<\min_{k_{b}\in\mathcal{K}_{\textsc{benign}}}\gamma(q(\mathcal{T}),k_{b})$
When the client processes $\mathcal{T}\in\mathcal{D}$, it computes $q(\mathcal{T})$ and selects the top-$N$ keys with the smallest distance $\gamma$ (Eq. 1). Since all $N$ adversarial keys have a smaller distance to $q(\mathcal{T})$ than all MNM-N benign keys, the top-$N$ selected keys must be exactly the set $\mathcal{K}_{\textsc{adv}}$.
The adversary’s guessing rule $\mathcal{A}_{\mathsf{GUESS}}$ predicts $b^{\prime}=1$ if all adversarial prompts $\mathcal{P}_{\textsc{adv}}$ are updated. Since $b=1$, these prompts will be selected and updated. Therefore, $\Pr[b^{\prime}=1\mid b=1]=1$. ∎
Lemma 3.2.
Let $q(x)\sim\mathcal{N}(k_{b_{i}},\sigma_{i}^{2}I)$ be a non-member query from benign cluster $i$. The probability $\Pr(E_{i})$, that $q(x)$ selects all $N$ adversarial keys $\mathcal{K}_{ADV}=\{k_{a_{1}},\dots,k_{a_{N}}\}$ as its $N$ closest centroids, is bounded by:
$\Pr(E_{i})\leq\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(\frac{(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}}{\sigma_{i}\|k_{a_{j}}-k_{b_{l}}\|}\right)$
where $\Phi(\cdot)$ is the Cumulative Distribution Function (CDF) of the standard normal distribution.
Proof.
Let $E_{i}$ be the event that a query $q(x)$, drawn from the $i$-th benign cluster (i.e., $q(x)\sim\mathcal{N}(k_{b_{i}},\sigma_{i}^{2}I)$), selects all $N$ adversarial keys. This occurs when all adversarial keys have a higher similarity to $q(x)$ than all MNM-N benign keys. This is formally defined as:
$E_{i}=\left\{\min_{j=1...N}q(x)^{T}k_{a_{j}}>\max_{l=1...M-N}q(x)^{T}k_{b_{l}}\right\}$
Here, we omit the normalization factors, since both the query and all keys are $\ell_{2}$-normalized. This is an intersection of $N\times(M-N)$ pairwise events, $E_{i}=\bigcap_{j=1}^{N}\bigcap_{l=1}^{M-N}A_{jl}$, where:
$A_{jl}=\left\{q(x)^{T}k_{a_{j}}>q(x)^{T}k_{b_{l}}\right\}\iff A_{jl}=\left\{(k_{a_{j}}-k_{b_{l}})^{T}q(x)>0\right\}$
The probability of an intersection of events is less than or equal to the probability of the single least likely event in that set. This gives us a strict upper bound:
$\Pr(E_{i})=\Pr\left(\bigcap_{j,l}A_{jl}\right)\leq\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Pr(A_{jl})$
Now, we find the probability of a single pairwise event $A_{jl}$. Let $Y_{jl}=(k_{a_{j}}-k_{b_{l}})^{T}q(x)$. Based on Assumption 2, $q(x)$ is a multivariate Gaussian $q(x)\sim\mathcal{N}(k_{b_{i}},\sigma_{i}^{2}I)$. The variable $Y_{jl}$ is a linear projection of $q(x)$, so it also follows a 1D Gaussian distribution. We find its mean $E[Y_{jl}]$ and variance $\text{Var}[Y_{jl}]$:
Mean:
$E[Y_{jl}]$ $=E[(k_{a_{j}}-k_{b_{l}})^{T}q(x)]=(k_{a_{j}}-k_{b_{l}})^{T}E[q(x)]=(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}$
Variance:
$\text{Var}[Y_{jl}]$ $=\text{Var}((k_{a_{j}}-k_{b_{l}})^{T}q(x))=(k_{a_{j}}-k_{b_{l}})^{T}\text{Cov}(q(x))(k_{a_{j}}-k_{b_{l}})$
$=(k_{a_{j}}-k_{b_{l}})^{T}(\sigma_{i}^{2}I)(k_{a_{j}}-k_{b_{l}})=\sigma_{i}^{2}\|k_{a_{j}}-k_{b_{l}}\|^{2}$
We want to find $\Pr(A_{jl})=\Pr(Y_{jl}>0)$. We standardize $Y_{jl}$ to $Z\sim\mathcal{N}(0,1)$ as:
$\Pr(Y_{jl}>0)=\Pr\left(Z>\frac{0-E[Y_{jl}]}{\sqrt{\text{Var}(Y_{jl})}}\right)=\Pr\left(Z>-\frac{E[Y_{jl}]}{\sqrt{\text{Var}(Y_{jl})}}\right)$
Using the identity $\Pr(Z>-z)=\Phi(z)$, the probability is $\Phi(z_{ijl}(\mathcal{K}_{ADV}))$, where the $z$-score is:
$z_{ijl}(\mathcal{K}_{ADV})=\frac{E[Y_{jl}]}{\sqrt{\text{Var}(Y_{jl})}}=\frac{(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}}{\sigma_{i}\|k_{a_{j}}-k_{b_{l}}\|}$
Substituting this result into the original inequality, we have the final bound:
$\Pr(E_{i})\leq\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(\frac{(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}}{\sigma_{i}\|k_{a_{j}}-k_{b_{l}}\|}\right)$
Theorem 3.3 (False Positive Rate).
The per-sample False Positive Rate (FPR) is bounded by:
$FPR=\Pr[b^{\prime}=1\mid b=0]\leq\max_{1\leq i\leq M-N}\left(\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(z_{ijl}\right)\right)$
where $\mathcal{K}_{ADV}=\{k_{a_{1}},\dots,k_{a_{N}}\}$ is the set of $N$ adversarial keys drawn from the shell $S$, and $z_{ijl}$ is the $z$-score:
$z_{ijl}=\frac{(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}}{\sigma_{i}\|k_{a_{j}}-k_{b_{l}}\|}$
Proof.
With batch size of $1$, the client has a single non-member sample $x$. The FPR is the probability of the event $E_{FP}$ that $q(x)$ selects all $N$ adversarial keys.
$FPR=\Pr(E_{FP})=\Pr\left(\min_{j=1...N}q(x)^{T}k_{a_{j}}>\max_{l=1...M-N}q(x)^{T}k_{b_{l}}\right)$
Let $p_{i}$ be the prior probability that a single non-member sample $x$ belongs to benign cluster $i$ (such that $\sum_{i=1}^{M-N}p_{i}=1$). We find this probability by summing over all MNM-N benign clusters that $q(x)$ could be drawn from:
$FPR=\sum_{i=1}^{M-N}\Pr(E_{FP}\mid q(x)\in\text{cluster }i)\cdot p_{i}$
where $p_{i}=\Pr(q(x)\in\text{cluster }i)$ is the prior probability for cluster $i$.
The term $\Pr(E_{FP}\mid q(x)\in\text{cluster }i)$ is exactly the probability $P(E_{i})$ in Lemma 3.2. We substitute its bound:
$\Pr(E_{FP}\mid i)\leq\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(z_{ijl}(\mathcal{K}_{ADV})\right)$
We place this bound into the above sum:
$FPR\leq\sum_{i=1}^{M-N}p_{i}\cdot\left(\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(z_{ijl}(\mathcal{K}_{ADV})\right)\right)\leq\max_{1\leq i\leq M-N}\left(\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(z_{ijl}(\mathcal{K}_{ADV})\right)\right)$
A.8 Advantage and Attack Success Rate of Individual models
Figures 1012 present a detailed comparison between our proposed PromptMIA and a naive membership inference baseline across three backbone models (ViT-B/32, ConViT, and DeiT) on all four datasets (CIFAR10, CIFAR100, TinyImageNet, and FourDataset). Each subplot reports both the Advantage and the Attack Success Rate (ASR) as a function of the client batch size. PromptMIA performs worse on FourDataset compared to other dataset. This is also the dataset with the lowest predictive accuracy under no DP (see Table 2). Results are averaged across 5 runs.
Figure 10: PromptMIA vs Naive attack results against ViT-B/32. Each subplot shows Advantage and Attack Success rate w.r.t Batch Size across CIFAR10, CIFAR100, TinyImageNet, and FourDataset.
Figure 11: PromptMIA vs Naive attack results against ConViT. Each subplot shows Advantage and Attack Success rate w.r.t Batch Size across CIFAR10, CIFAR100, TinyImageNet, and FourDataset.
Figure 12: PromptMIA vs Naive attack results against Deit. Each subplot shows Advantage and Attack Success rate w.r.t Batch Size across CIFAR10, CIFAR100, TinyImageNet, and FourDataset.
A.9 Detailed results on outlier detection
Table 3 reports the full precision, recall, and F1 scores of classical anomaly detection methods applied to the task of detecting adversarial keys in the global prompt pool across all datasets and backbone models. We observe that IsolationForest achieves the highest recall ($\approx 1.0$) on every setting, meaning it successfully flags almost all injected adversarial keys. However, its precision is low (typically 0.180.180.370.37), indicating that many benign keys are incorrectly labeled as adversarial, leading to high false positives. OneClassSVM shows slightly better precision ($\sim 0.15$$0.30$) but still suffers from moderate recall and poor F1 scores. LocalOutlierFactor and EllipticEnvelope fail almost entirely in this scenario, often yielding zero detection or near-zero scores. Visualization of these outlier detecion methods are given in Fig. 5. Moreover, these methods consistently misclassify benign keys as adversarial even when no adversarial keys are present ( see Fig. 13).
(a) Isolation Forest
(b) Local Outlier Factor
(c) OneClassSVM
(d) EllipticEnvelope
Figure 13: Visualization of outlier detection methods on CIFAR-10 trained ViT-B32. Blue keys are benign keys. Red keys are adversarial keys. Crossed keys are flagged as outliers from the corresponding algorithm. Outlier detection methods still falsely flag benign keys as outliers when no adversarial keys are present.
Dataset Model Method Precision Recall F1
CIFAR10 ViT IsolationForest 0.2828 1.0000 0.4409
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.3052 0.4773 0.3723
EllipticEnvelope 0.0000 0.0000 0.0000
ConViT IsolationForest 0.2330 1.0000 0.3779
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.2977 0.4894 0.3702
EllipticEnvelope 0.0079 0.0213 0.0116
DeiT IsolationForest 0.2746 1.0000 0.4308
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.1734 0.4602 0.2519
EllipticEnvelope 0.0000 0.0000 0.0000
CIFAR100 ViT IsolationForest 0.2576 1.0000 0.4096
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.2218 0.5196 0.3109
EllipticEnvelope 0.0000 0.0000 0.0000
DeiT IsolationForest 0.3623 1.0000 0.5319
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.2257 0.5450 0.3192
EllipticEnvelope 0.0022 0.0050 0.0030
ConViT IsolationForest 0.3128 1.0000 0.4766
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.1592 0.5045 0.2420
EllipticEnvelope 0.0000 0.0000 0.0000
TinyImageNet ViT IsolationForest 0.1794 1.0000 0.3042
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.1540 0.4938 0.2348
EllipticEnvelope 0.0000 0.0000 0.0000
DeiT IsolationForest 0.1444 1.0000 0.2524
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.2132 0.5765 0.3113
EllipticEnvelope 0.0000 0.0000 0.0000
ConViT IsolationForest 0.3014 1.0000 0.4632
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.2561 0.4873 0.3358
EllipticEnvelope 0.0000 0.0000 0.0000
FourDataset ViT IsolationForest 0.1944 1.0000 0.3255
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.1485 0.5000 0.2290
EllipticEnvelope 0.0087 0.0167 0.0114
DeiT IsolationForest 0.3743 1.0000 0.5447
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.1412 0.5000 0.2202
EllipticEnvelope 0.0045 0.0104 0.0062
ConViT IsolationForest 0.2896 1.0000 0.4492
LocalOutlierFactor 0.0000 0.0000 0.0000
OneClassSVM 0.1498 0.4387 0.2233
EllipticEnvelope 0.0054 0.0094 0.0069
Table 3: Precision, Recall, and F1 of Outlier Detection across datasets, models, and methods.
A.10 Hyperparameter Analysis
To isolate the effect of each hyperparamter, all experiments in this section are conducted using Vit-B32 model and CIFAR100 dataset.
Global Prompt Pool size $M$: Increasing the pool size strengthens the attack. Larger $M$ (e.g., $M=24$) yields consistently higher attack success rates across all batch sizes, while smaller pools (e.g., $M=12$) slightly weaken the attack as batch size grows. When the pool size increases, the probability of the adversarial keys being selected when $\mathcal{T}\notin\mathcal{D}$ decreases, increasing FPR. Since the server is in control of the training protocol and the global prompt pool, they can choose the value $M$. See Fig. 14(a).
Prompt selection size $N$: Increasing $N$ slightly weakens the attack. Again, since the server controls the training protocol, they can also most likely dictate the choice of $N$. See Fig. 14(b).
Impact of $\delta_{\min}$: Without any defense, increasing $\delta_{\min}$ reduces the attack success rate; therefore one may choose $\delta_{\min}$ arbitrarily small (e.g. $\delta_{\min}=0.02$). However, when the client employs input-noise perturbation, the adversary benefits from a larger $\delta_{\min}$ (0.20.3)(0.2-0.3), yet not so large that all adversarial keys collapse onto the target query. See Fig. 14(c).
Impact of $\Delta$: Empirically, increasing $\Delta$ reduces inference accuracy, and a relatively small $\Delta$ does not make the attack detectable by traditional anomaly-detection methods. Therefore we choose $\Delta$ to be modest (e.g. $\Delta=0.05$). However, $\Delta$ should not be so small that the adversarial keys become indistinguishably close to one another. See Fig. 14(d).
Impact of $\beta$: Increasing $\beta$ increase alignment between adversarial and benign keys (see Fig. 9), but at the cost reducing attack success rate. We find $\beta=0$ to be sufficient against traditional outlier detection methods, however carefully tuning $\beta$ might be helpful against a potentially more potent anomaly detection algorithm. See Fig. 14(e).
Impact of number of training rounds: PromptMIA much higher ASR against models that have been trained for a few rounds compared to randomly initialized keys. This is reflected in our theorectical findings in Section 3.2.3 that FPR is lower when the benign data forms tight and compact clusters in the query space around the benign keys, which happens naturally during the training process. See Fig. 14(f) and Fig. 15.
(a) Impact of prompt pool size $M$
(b) Impact of prompt selection size NN
(c) Impact of $\delta_{\min}$
(d) Impact of $\Delta$
(e) Impact of $\beta$
(f) Impact of number of training rounds
Figure 14: Ablation study on PromptMIA. Each subfigure shows the effect of one parameter: (a) $M$, (b) NN, (c) $\delta_{\min}$, (d) $\Delta$, (e) $\beta$, and (f) training rounds.
A.11 Training Dynamics of Benign Keys
In federated prompt tuning, the keys in the global prompt pool gradually adapt to represent the distribution of clients’ data. Early in training (Round 0), benign keys are randomly initialized and do not reflect the query feature space. As training proceeds, the selected keys are pulled toward their associated query vectors, causing the benign keys to migrate toward dense regions of the feature space. Over multiple communication rounds, these keys stabilize and effectively act as cluster centroids for groups of similar queries. Figure 15 visualizes this process, showing how random keys become structured and aligned with the data distribution after the training process (Fig. 15).
(a) Round 0 (Random weight)
(b) Round 1
(c) Round 20
(d) Round 60
Figure 15: Visualization of distribution of benign keys (blue) and query vectors $q(x)$ (green) across training rounds.
A.12 Membership Inference under very large batch sizes
To assess whether further increasing the batch size can mitigate membership leakage, we conducted an additional set of experiments using an extreme configuration with batch size set to 1024. As seen in Fig. 16, the attack success rate remains close to 80% on FourDataset and more than 80% on others. We also note that using such batch size is often not possible in practical scenarios where low-resource edge devices cannot afford high VRAM consumption. Defense using large batch size is therefore ineffective and impractical.
Figure 16: Attack success rate of PromptMIA under increasing batch size ( up to 1024) across different datasets. The results consistently show that the attack success rate remains high even under extremely large batch size.
A.13 Performance of PromptMIA against text and multimodal data.
To demonstrate that PromptMIA extends beyond Vision Transformer, we conduct additional experiments on the UPMC Food-101 dataset (Gallo et al., 2020) which is a multimodal image-text benchmark containing image–caption pairs. For these experiments, we use the pretrained Vision-and-Language Transformer (ViLT) with a frozen image encoder and frozen LLM-based text encoder (i.e., BERT). We additionally evaluate a text-only configuration by providing only textual inputs. In both the multimodal and text-only cases, PromptMIA achieves strong attack success rates, confirming that the attack is not restricted to the vision domain (see Fig. 17).
Figure 17: Attack Success Rate of PROMPTMIA under multimodal and text input modality.
A.14 Performance under non-heterogenous settings
In addition to Fig. 3 in the main text, we have also run additional sensitivity studies on the attack success rate under more heterogeneous setting where such assumption might be less accurate. In particular, we adopt a Dirichlet-based heterogeneous data partitioning strategy. Under this setup, each client observes samples from all classes, but the class proportions differ across clients. We generate these non-IID splits by sampling class proportions for each client from a Dirichlet($\alpha\cdot\mathbf{1}_{s}$) distribution over an $s$-dimensional simplex, where $s$ is the number of classes and $\alpha$ is the concentration parameter with $\alpha=0.1$ and $\alpha=0.5$.
Figure 18: Attack success rate of PROMPTMIA against prompt-based FL on CIFAR-10 under different heterogeneity settings. Even on extremely large batch size, the attack success rate remains highly significant at more than 87%.
To validate Assumption 2, we visualize the distributions of benign keys and non-member queries for models trained on CIFAR-10 with Dirichlet parameters $\alpha=0.1$ and $\alpha=0.5$ below. Both plots in fact visualize empirical prompts clusters that resemble mixtures of Gaussian. Our experiments also show that the adversarial advantage and attack success rate of PromptMIA in the more extreme non-IID setting ($\alpha=0.1$) remains significant which supports our observation above on how Assumption 2 reasonably fits the empirical prompt clusters.
Figure 19: Visualization of prompt clusters produced by PFPT when trained on CIFAR10 with heterogeneous setting $\alpha=0.1$
Figure 20: Visualization of prompt clusters produced by PFPT when trained on CIFAR10 with heterogeneous setting $\alpha=0.5$